# The EnvDesign Model: A Method to Solve the Environment Design Problem

### v3

@@ 1.2 Pre-Production Testing Environments
This paper describes the evolution of this problem, formulates it as an $NP$-hard combinatorial constrained optimization problem called the “environment design” problem, and presents the EnvDesign model (designed and developed by the authors of this paper), which solves the environment design problem.
@@ 2.1 Evolution of the Problem
The clique cover problem is a $NP$-complete decision problem (CAI, ) (referred to as “partition into cliques” in the cited source). By definition, all $NP$-complete problems are also $NP$-hard (CLR4, ), so the clique cover problem is also $NP$-hard. Since the clique cover problem is $NP$-hard, it likely cannot be solved in polynomial time (Optimizn, ).
The minimum clique cover problem is an optimization problem, where the optimal solution is the clique cover consisting of the smallest number of cliques. Any optimization problem is at least as hard as its corresponding decision problem (CLR4, ), which in this case is the $NP$-hard clique cover problem. It follows that the minimum clique cover problem, and the more constrained version of it presented here, are also $NP$-hard and the optimal solution likely cannot be obtained in polynomial time.
@@
In practice, we often have more nodes available for testing than are required to satisfy the coverage constraint. Given $n$ nodes for testing, it is best to use all $n$ nodes and perform as much testing as possible.
The configurations for these $n$ nodes must be determined in such a way that the coverage constraint is satisfied and, to the greatest extent possible, the desired/target distribution of dimension values is matched as closely as possible in the collection of testing configurations (schedule). Let this closeness between true and target distributions be quantified by an objective function $O(S)$, where $S$ is a schedule. When $O(S)=0$, the distribution matches exactly with the target distributions.
For some testing scenarios, only some of the dimension values in a dimension need to be covered and other dimension values need to be excluded. These constraints are represented by a scope $C=(I=I_{1}\cup I_{2}\cup I_{3},X=X_{1}\cup X_{2}\cup X_{3})$, where $I$ contains the include scope (dimension values in each dimension that need to be covered) and $X$ contains the exclude scope (dimension values in each dimension that need to be excluded). We assume that the two sets don’t intersect for any dimension (if they did, the problem would be immediately infeasible since how can a dimension value be included and excluded at the same time).
Since the environment design problem can be reduced to the $NP$-hard clique cover problem and seeks to minimize the value of an objective function, it is a $NP$-hard constrained optimization problem.
@@ 2.2 Formulation of the Environment Design Problem
    The number of testing configurations $n\in\mathbb{Z}^{+}$.
@@
    Scope $C=(I=I_{1}\cup\dots\cup I_{d},X=X_{1}\cup\dots\cup X_{d})$, $\forall i,I_{i}\cap X_{i}=\emptyset$, which consists of collections of values to include ($I$) and exclude ($X$) for each dimension.
@@
    Each testing configuration has $d$ dimension values, $\forall s\in S,|s|=d$
@@ 5 Conclusion
This paper discusses the evolution of the environment design problem and formulates it as an $NP$-hard constrained optimization problem, where the goal is to dynamically provide a diverse testing schedule tailored to current testing priorities.

### v4

@@ 1.2 Pre-Production Testing Environments
This paper describes the evolution of this problem, proves that it is $NP$-hard, formulates it as a combinatorial constrained optimization problem called the “environment design” problem, and presents the EnvDesign model (designed and developed by the authors of this paper), which solves the environment design problem.
@@ 2.1 Evolution of the Problem
The clique cover problem is an $NP$-complete decision problem that asks if there exists a collection of cliques, the size of which is no greater than a given bound, that covers all the vertices in a given graph (Karp, ). By definition, all $NP$-complete problems are also $NP$-hard (CLR4, ), so the clique cover problem is also $NP$-hard. Since the clique cover problem is $NP$-hard, it likely cannot be solved in polynomial time (Optimizn, ).
The minimum clique cover problem is an optimization problem, where the optimal solution is the clique cover consisting of the smallest number of cliques. Any optimization problem is at least as hard as its corresponding decision problem (CLR4, ), which in this case is the $NP$-hard clique cover problem. It follows that the minimum clique cover problem is also $NP$-hard and the optimal solution likely cannot be obtained in polynomial time.
Hence, the environment design problem has now evolved into a more constrained version of the $NP$-hard minimum clique cover problem.
@@
In practice, we often have a specified number of nodes for testing, $n$. It follows that the size of the clique cover must be less than or equal to $n$. Now that the size of the clique cover is bounded, the environment design problem has become a more constrained version of the $NP$-hard clique cover problem (the decision problem with a bound, as opposed to the minimum clique cover problem, an optimization problem without a bound).
Although all $n$ nodes may not be needed to satisfy the coverage constraint, it is best to use all $n$ nodes and perform as much testing as possible. The configurations for these $n$ nodes must be determined in such a way that the coverage constraint is still satisfied and, to the greatest extent possible, the desired/target distribution of dimension values is matched as closely as possible in the collection of testing configurations (schedule). Let this closeness between true and target distributions be quantified by an objective function $O(S)$, where $S$ is a schedule. When $O(S)=0$, the distribution matches exactly with the target distributions.
For some testing scenarios, only some of the dimension values in a dimension need to be covered and other dimension values need to be excluded. These constraints are represented by a scope $C=(I=I_{1}\cup I_{2}\cup I_{3},X=X_{1}\cup X_{2}\cup X_{3})$, where $I$ contains the include scope (dimension values in each dimension that need to be covered) and $X$ contains the exclude scope (dimension values in each dimension that need to be excluded). We assume that the two sets don’t intersect for any dimension (if they did, the problem would be immediately infeasible since a dimension value cannot be included and excluded at the same time).
Since the environment design problem is a more constrained version of the $NP$-hard clique cover problem, is an $NP$-hard problem itself (proof provided in section 7.2 of this paper), and seeks to minimize the value of an objective function, it is an $NP$-hard constrained optimization problem.
@@ 2.2 Formulation of the Environment Design Problem
    The number of testing configurations $n\in\mathbb{Z}^{+}$
@@
    Scope $C=(I=I_{1}\cup\dots\cup I_{d},X=X_{1}\cup\dots\cup X_{d})$, $\forall i,I_{i}\cap X_{i}=\emptyset$, which consists of collections of dimension values (for each dimension) that must be included in ($I$) and excluded from ($X$) the output.
@@
    Each testing configuration has $d$ dimension values $\forall s\in S,|s|=d$, exactly one dimension value from each dimension $\forall s\in S,\forall i\in\{1,\dots,d\},|s\cap V_{i}|=1$
@@ 5 Conclusion
This paper discusses the evolution of the environment design problem, proves that it is $NP$-hard, and formulates it as a constrained optimization problem, where the goal is to dynamically provide a diverse testing schedule tailored to current testing priorities.